1.Determine three consecutive numbers such that the sum of twice the smallest number, three
times the middle number, plus the largest number is 53.
2.Mark is 6 years older than Stephanie. Two times Mark’s age is 8 less than 4 times Stephanie’s
age. What are Mark and Stephanie’s ages?

Answer by alanc(27)

1.)
Assign variables.
Let the 3 consecutive numbers be: a, b, c.
a < b < c
2a + 3b + c = 53
also consecutive means b = a + 1, c = a + 2
2a + 3*(a + 1) + (a + 2) = 53
2a + 3a + 3 + a + 2 = 53
6a + 5 = 53
6a = 48
a = 8,
The consecutive numbers are 8, 9, and 10
2.) M = 6 + S , 2M = 4S - 8
or M = 2S - 4, 2S - 4 = 6 + S, S = 6 + 4, S = 10
Then Mark is 6 + 10 = 16 years old
Stephanie is 10 years old.
Mark is 16 years old
3.)